A pit was 50 handbreadths in depth: A lion climbed up the pit 1/7 handbreadth every day and fell back 1/9 handbreadth. How long would it take him to get out of the pit?
Answer by fractalier(6550)

If he climbs up 1/7 and falls back 1/9 each day, he actually gains
1/7 - 1/9 = 2/63 handbreadth each day.
Now divide 50 by 2/63 and get
50 / (2/63) = 3150/2 = 1575 days
(You better be feeding him...)